OpenStudy (anonymous):

can someone please check my answer so the problem was solve for x 1/5x + 2/y = 1/x and my answer was x=5-10 im not sure if im right

jhonyy9 (jhonyy9):

solve for x : 1/5x + 2/y =1/x y+10x=5y 10x=4y x=4y/10 x=2y/5